# How Does A Text Preprocessing Pipeline Affect Syntactic Ontology Matching?

### v8

@@ Abstract.
The classical text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for syntactic ontology matching (OM). However, the lack of standardisation in text preprocessing creates diversity in mapping results. In this paper, we investigate the effect of the text preprocessing pipeline on syntactic OM in 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments. We find that Phase 1 text preprocessing (Tokenisation and Normalisation) is more effective than Phase 2 text preprocessing (Stop Words Removal and Stemming/Lemmatisation). We propose two novel approaches to repair unwanted false mappings caused by Phase 2 text preprocessing. One is an ad hoc logic-based repair approach that employs an ontology-specific check to find common words that cause false mappings. These words are stored in a reserved word set and applied before the text preprocessing. By leveraging the power of large language models (LLMs), we also propose a post hoc LLM-based repair approach. This approach utilises the strong background knowledge provided by LLMs to repair non-existent and counter-intuitive false mappings after the text preprocessing. It also overcomes the tendency towards unstable true mappings by injecting the classical text preprocessing pipeline via function calling. The experimental results show that these two approaches can improve the matching correctness and the overall matching performance.
Ontology Matching, Text Preprocessing
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 29th International Conference on Extending Database Technology; March 24–27, 2026; Tampere, Finlandisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Ontology engineering
1. Introduction
Ontology matching (OM) is essential for enabling interoperability between heterogeneous ontologies. An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences (Euzenat and Shvaiko, 2013). A classical OM system usually contains syntactic matching, lexical matching, and semantic matching. This multilayer architecture has been implemented in several successful OM systems, such as LogMap (Jiménez-Ruiz and Cuenca Grau, 2011; Jiménez-Ruiz et al., 2011), AgreementMakerLight (AML) (Faria et al., 2013, 2014), and FCA-Map (Zhao et al., 2018; Li et al., 2021).
Syntactic matching captures “anchor mappings” and provides the foundation for the latter lexical and semantic matching. There are many strategies for extracting syntactic information from an ontology entity. These include older approaches Bag-of-Words (e.g. TF-IDF (Sammut and Webb, 2010)), popular word embedding models (e.g. Word2Vec (Mikolov et al., 2013)), and state-of-the-art language models (e.g. BERT (Devlin et al., 2019)). Despite the diversity of the strategies used, they all apply the text preprocessing pipeline to clean the text data before using the model.
@@ 1. Introduction
The classical text preprocessing pipeline consists of a set of steps to segment, reconstruct, analyse, and process the information in the text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) (Manning et al., 2008). T is the process of breaking the text into the smallest units (i.e. tokens). N is the process of transforming these different tokens into a single canonical form. R is the process of removing filler words that usually carry little meaning and can be omitted in most cases. S/L is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word. Table 1 shows examples of using the text preprocessing pipeline to find mappings in syntactic OM. In the examples, we use whitespace to split the tokens, normalise all the words into lowercase, remove the common stop words “has”, “a” and “is”, and apply stemming/lemmatisation to find the simplest word form “review” from “reviews” and “reviewing”.
No. Entity Name Matching Output Preprocessing 1 ProgramCommittee Program Committee T Program_Committee 2 ProgramCommittee program committee T + N Program_committee 3 has_a_review review T + N + R hasReview 4 reviews review T + N + R + S/L isReviewing
Table 1. Examples of using the classical text preprocessing pipeline in syntactic ontology matching: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
@@
A number of OM systems use the text preprocessing pipeline for syntactic OM, but few studies question its role in optimising the matching performance. Our study fills this gap. We first conduct a comprehensive experimental analysis of the text preprocessing pipeline used in syntactic OM across a wide range of domains, aiming to explain the behaviour of the text preprocessing pipeline. In each step, text preprocessing methods are evaluated for correctness and completeness. While several steps of the text preprocessing pipeline are currently less effective, we propose two novel pipeline repair approaches. These two methods utilise the logic of ontological representations and the power of large language models (LLMs), showing promising results for repairing weaknesses of the text preprocessing pipeline. Specifically, this paper makes the following contributions:
@@
  • We compare the performance of the text preprocessing pipeline on 8 datasets of the Ontology Alignment Evaluation Initiative (OAEI) (OAEI Community, [n. d.]) with 49 distinct alignments. While there is a significant improvement using Tokenisation (T) and Normalisation (N) in syntactic OM, Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective. We find common inappropriate stop words removal, over-stemming, and over-lemmatisation.
  • We propose two approaches to repair the less effective text preprocessing pipeline in syntactic OM: (1) ad hoc logic-based repair and (2) post hoc LLM-based repair. Experimental results show that these two methods can improve the matching correctness and the overall matching performance of syntactic OM.
@@
The remainder of the paper is organised as follows. Section 2 and 3 review the related work and the preliminaries of OM and syntactic matching. Section 4 analyses the text preprocessing pipeline used in syntactic OM. Section 5 proposes text preprocessing pipeline repair approaches and experimentally validates their performance. Section 7 highlights the contributions of this work and concludes the paper.
2. Related Work
Syntactic OM considers only the meaning of an entity’s name or label, ignoring its lexical and structural context in an ontology. Correct syntactic matches are often implicit and usually require extra human observation and domain knowledge (Euzenat and Shvaiko, 2013). Text preprocessing is introduced to automate this process. The use of text preprocessing in syntactic OM can be traced back to the early stages of OM systems, which were initially developed and widely used as a basic component to generate linguistic-based mappings. AgreementMaker (Cruz et al., 2009) has a normaliser to unify the textual information of the entities. SAMBO (Lambrix and Tan, 2006) uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes. In RiMOM (Li et al., 2009), the context information of each entity is viewed as a document. The text in each document is preprocessed with tokenisation, stop words removal, and stemming. Machine learning (ML) models have emerged for modern OM systems. While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model. For example, BERTMap (He et al., 2022) uses BERT’s inherent WordPiece tokeniser to build the subword of each entity. DeepOnto (He et al., 2024) extends normalisation to axioms using EL embedding models (Kulmanov et al., 2019). The ML extension of LogMap (Chen et al., 2021) reuses the seed mappings of traditional systems, where each entity is split into its component words and the mapping is based on pairwise component similarities. However, to the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have evaluated the effect of the entire text preprocessing pipeline on syntactic OM.
Early explorations in (He et al., 2023; Norouzi et al., 2023) show the potential of using LLMs for OM, but matching by LLM prompting alone commonly results in poor performance. OLaLa (Hertling and Paulheim, 2023) proposes a prototype and states that the LLMs for OM include a number of decisions on model and parameter combinations, and there is no one-model-fits-all solution. Works on both Agent-OM (Qiang et al., 2023) and LLM4OM (Giglou et al., 2024) point out that LLMs have limited context length and inherited hallucinations. While LLM4OM highlights the importance of integrating the Retrieval-Augmented Generation (RAG) (Lewis et al., 2020), Agent-OM introduces an agent-based framework and enhances the compatibility of LLMs via function calling (OpenAI, 2023) for OM to improve matching performance. The authors in (Taboada et al., 2025) state that the use of LLMs for OM has a high computational overhead and apply a prioritised depth-first search (PDFS) strategy on the retrieve-then-prompt pipeline to optimise candidate selection. Recent works in (Amini et al., 2024; Zamazal, 2024; Sousa et al., 2025) also utilise LLMs and LLM embeddings for complex OM. However, these works focus on building a sophisticated framework for using LLMs for OM, and less attention has been paid to the subtask of using LLMs for syntactic matching. For syntactic OM, the challenges of using LLMs are twofold. On one hand, LLM hallucinations may produce incorrect “seed mappings”. Such mappings are toxic for the subsequent lexical and semantic matching. On the other hand, it is difficult to select a generic LLM prompting strategy from a number of LLM prompting templates and strategies.
3. Preliminaries
@@
An ontology provides a formalised conceptualisation of domain knowledge (Gruber, 1993). The basic component of an ontology is the entity. Each entity is assigned a unique identifier (i.e. IRI) and represents a unique concept of the domain. Entities can be a concept (i.e. class), relation (i.e. property), or instance (i.e. individual) (Noy et al., 2001). Ideally, the entity names are not duplicated within a single ontology to minimise confusion and overlap. It is recommended that ontologies representing similar domain knowledge use similar terminologies for entity names (OBO Foundry, [n. d.]).
@@
Ontology matching (OM) aims to capture and align similar concepts that occur in different ontologies. Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), OM establishes mappings between pairs of entities drawn from each of the two ontologies. A correspondence (a.k.a. a mapping) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence of the correspondence (or may be absent). The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships (Euzenat and Shvaiko, 2013). OM is a complex process that combines syntactic matching, lexical matching, and semantic matching. Each level of matching may use different matching techniques, for example, text preprocessing for syntactic matching, text vectorisation for lexical matching, and logical reasoning for semantic matching. This study concentrates on the common practice of using the text preprocessing pipeline in syntactic OM. We focus only on the equivalence relationship ($\equiv$) and evaluate the effect of text preprocessing on syntactic matching to produce the “anchor mappings” on which to base any subsequent lexical and semantic matching. We address only equivalence because subsumption is typically dealt with in a later semantic and structural matching phase of OM.
4. Text Preprocessing Pipeline Analysis
In this section, we discuss the role of the text preprocessing pipeline in syntactic OM and report our experiments that study the effect of different text preprocessing options on the OAEI datasets.
4.1. Experiment Setup
@@
Figure 1 shows the experiment setup to analyse the text preprocessing pipeline. Alignment (A) is generated via the text preprocessing pipeline. For both $O_{s}$ and $O_{t}$, we retrieve the entities from classes (owl:Class) and properties (both owl:ObjectProperty and owl:DatatypeProperty). For those ontologies where the names of the concepts are not textual (e.g. a code or numerical identifier), we retrieve the meaningful text from entity labels (rdfs:label or other annotation properties) instead. Then, we apply the text preprocessing pipeline method $f(\cdot)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file. While OM methods will often follow text preprocessing with a more sophisticated strategy to determine equivalence of entities, we define equivalence to mean identical text generated by preprocessing alone. This enables us to directly evaluate the preprocessing performance as a standalone contribution to OM. We compare the generated alignment (A) with the reference (R) to evaluate the performance of the text preprocessing pipeline on syntactic OM.
@@
Figure 1. Experiment setup to analyse the text preprocessing pipeline: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
[…]
Chart of experiment setup to analyse the text preprocessing pipeline.
  1. (1)
@@
    Selected OAEI tracks: We select 8 OAEI tracks stored in the Ontology Matching Evaluation Toolkit (MELT) (Hertling et al., 2019) public repository (retrieved January 1, 2025). Table 2 shows the details of the selected tracks. Each track may contain more than one alignment corresponding to different pairs of ontologies. For example, the Largebio Track has multiple reference files, pairing FMA (Rosse and Mejino, 2003) and NCI (Golbeck et al., 2003), FMA and SNOMED (Donnelly, 2006), and SNOMED and NCI, respectively. The number of alignments for each track is given in the table. There are 49 distinct alignments evaluated in this study. Figure 2 shows the number of entities in each alignment.
    Table 2. Selected OAEI tracks and the number of alignments.
@@
    Name Domain Alignments Anatomy Human and Mouse Anatomy 1 Biodiv Biodiversity and Ecology 9 CommonKG Common Knowledge Graphs 3 Conference Conference 24 Food Food Nutritional Composition 1 Largebio Biomedical 6 MSE Materials Science & Engineering 3 Phenotype Disease and Phenotype 2
@@
    Figure 2. Number of entities in each alignment.
[…]
    Chart of number of entities in each alignment.
@@
    Compound words are frequently used in ontology naming conventions. For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO (Suchanek et al., 2007) or the Snake case “art_gallery” used in Wikidata (Vrandečić and Krötzsch, 2014). Figure 3(b) shows the frequency distribution of compound words in the selected OAEI tracks. For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y. However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”. Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
@@
    Figure 3. Frequency distribution of compound words. We exclude entities with more than 15 compound words because their proportion is less than 1%.
[…]
    Chart of frequency distribution of compound words.
  2. (2)
@@
    Selected pipeline methods: There are a variety of methods that can be used in a general text preprocessing pipeline, but not all of them are applicable to syntactic OM. We select the following methods in each text preprocessing step: (a) Tokenisation (T): We use full word tokenisation to separate the compound words. Experiments on several subword tokenisation methods (e.g. Byte Pair Encoding (Gage, 1994; Sennrich et al., 2016), WordPiece (Schuster and Nakajima, 2012; Devlin et al., 2019), and Unigram (Kudo, 2018; Kudo and Richardson, 2018)) show the results after de-tokenisation (i.e. converting the tokens to a sentence) are generally the same as full word tokenisation. We do not use sentence tokenisation (a.k.a. segmentation) because text retrieved from the entity’s name or label is commonly short. (b) Normalisation (N): includes lowercasing, HTML tag removal, separator formatting, and punctuation removal. Other methods that may potentially change the word semantics are excluded (e.g. removal of special characters and numbers). (c) Stop Words Removal (R): includes the most common English stop words defined in the Natural Language Toolkit (NLTK) (Bird et al., 2008). (d) Stemming/Lemmatisation (S/L): Stemming methods include Porter, Snowball, and Lancaster. Lemmatisation uses the NLTK Lemmatiser based on WordNet (Miller, 1995), and the word categorisation uses POS Tagging.
  3. (3)
@@
    Selected measures: There are four primitive measures in information retrieval: true positive (TP), false positive (FP), false negative (FN), and true negative (TN). In the context of OM, the evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R) validated by domain experts. Figure 4 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, $FP=A-R$, $FN=R-A$, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences. Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN. In the context of OM, since $C\times C^{\prime}$ is extremely large (the Cartesian product of $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$), TN is often much larger than TP, FP, and FN. This means that Accuracy (Acc) and Specificity (Spec) are close to 1, and they have no statistically significant difference across different alignments. We note that Precision (Prec) and Recall (Rec) contribute equally to $F_{\beta}$. Therefore, we choose Precision (Prec), Recall (Rec), and F1 Score ($\beta=1$) in this study. They are defined as follows.
    Figure 4. OM evaluation (Euzenat, 2007).
[…]
    Chart of OM evaluation (Euzenat, 2007).

### v10

# How Does A Text Preprocessing Pipeline Affect Ontology Matching?

@@ Abstract
The classical text preprocessing pipeline, comprising Tokenisation, Normalisation, Stop Words Removal, and Stemming/Lemmatisation, has been implemented in many systems for ontology matching (OM). However, the lack of standardisation in text preprocessing creates diversity in the mapping results. In this paper, we investigate the effect of the text preprocessing pipeline on 8 Ontology Alignment Evaluation Initiative (OAEI) tracks with 49 distinct alignments. We find that Tokenisation and Normalisation (categorised as Phase 1 text preprocessing) are more effective than Stop Words Removal and Stemming/Lemmatisation (categorised as Phase 2 text preprocessing). We propose two novel approaches to repair unwanted false mappings that occur in Phase 2 text preprocessing. One is an ad hoc logic-based repair approach used before text preprocessing, employing an ontology-specific check to find common words that cause false mappings. The other repair approach is the post hoc large language model (LLM)-based approach, used after text preprocessing, which utilises the strong background knowledge provided by LLMs to repair non-existent and counter-intuitive false mappings. The experimental results indicate that these two approaches can significantly improve the matching correctness and the overall matching performance.
I Introduction
@@
Ontology matching (OM) is crucial for enabling interoperability between heterogeneous ontologies, thereby facilitating the integration of disparate knowledge graphs (KGs) [1]. An OM process usually takes two ontologies as input, discovers mappings between entities, and produces a set of correspondences [2]. The classical text preprocessing pipeline is commonly used in OM. The pipeline consists of a set of steps to segment, reconstruct, analyse, and process text, namely Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L) [3]. T is the process of breaking the text into the smallest units (i.e. tokens). N transforms these different tokens into a single canonical form. R is the process of removing filler words that usually carry little meaning and can be omitted in most cases. S/L is used to deal with the grammatical variation of words, applying rules to find the simplest common form of the word.
Fig. 1 shows that two different ontology entities “reviews” and “isReviewing” can be matched via the text preprocessing pipeline as follows. After tokenisation, our tokens comprise a sequence of word units which we present as separated by white spaces within a token (“reviews” cf “is Reviewing’). We then normalise all words to lowercase (“reviews” cf “is reviewing”), remove the common stop words “is” (“reviews” cf “reviewing”), and finally apply stemming/lemmatisation to find the simplest word form “review” from both “reviews” and “reviewing” (“review” cf “review”).
Figure 1: An Example of using the text preprocessing pipeline in OM.
@@
A number of OM systems utilise the text preprocessing pipeline, but few studies have justified the choice of a specific text preprocessing method and systematically studied its role in optimising the matching performance. Our study fills this gap. We first conduct a comprehensive experimental analysis of the text preprocessing pipeline used in OM across a wide range of domains, aiming to explain the impact on OM of each step in the pipeline. In each step, text preprocessing methods are evaluated for correctness and completeness. We find that some steps of the text preprocessing pipeline are currently less effective. Based on this insight, we propose two novel pipeline repair approaches. These two methods utilise the logic of ontological representations and the power of large language models (LLMs), showing promising results for repairing weaknesses of the text preprocessing pipeline. Specifically, this paper makes the following contributions.
@@
  • We comprehensively evaluate the performance of each step in the text preprocessing pipeline on 8 datasets of the Ontology Alignment Evaluation Initiative (OAEI) [4] with 49 distinct alignments. While there is a significant improvement using Tokenisation (T) and Normalisation (N), Stop Words Removal (R) and Stemming/Lemmatisation (S/L) are currently less effective. We find the issue is caused by inappropriate stop words removal, over-stemming, and over-lemmatisation. We also compare the performance of (1) Stemming and Lemmatisation, (2) different stemming methods including Porter, Snowball, and Lancaster, and (3) Lemmatisation with and without Part-of-Speech (POS) Tagging.
  • We propose two approaches to repair the less effective text preprocessing pipeline: (1) ad hoc logic-based repair and (2) post hoc LLM-based repair. To reduce the number of false mappings, the logic-based repair approach applies an ad hoc check based on ontology-specific settings before text preprocessing, while the LLM-based repair utilises LLMs as a post hoc validator after text preprocessing. We examine the performance of OM across different LLMs and prompt templates. We find the optimal pipeline to combine the traditional text processing pipeline with modern LLMs for OM. The experimental results indicate that these two repair methods can significantly improve the matching correctness and the overall matching performance.
@@
The paper is organised as follows. Sections II and III review the preliminaries and related work of OM and the text preprocessing pipeline. Section IV analyses the text preprocessing pipeline used in OM. Section V proposes text preprocessing pipeline repair approaches and experimentally validates their performance. Section VI discusses the limitations of this work. Section VII concludes the paper.
II Preliminaries
@@
An ontology provides a formalised conceptualisation of domain knowledge [5]. The basic component of an ontology is the entity. Each entity is assigned a unique identifier (i.e. IRI) and represents a unique concept of the domain. Entities can be a concept (i.e. class), relation (i.e. property), or instance (i.e. individual) [6]. Ideally, the entity names are not duplicated within a single ontology to minimise confusion and overlap. It is recommended that ontologies representing similar domain knowledge use similar terminologies for entity names [7].
@@
Ontology matching (OM) aims to capture and align similar concepts that occur in different ontologies. Given a source ontology ($O_{s}$) and a target ontology ($O_{t}$), OM establishes mappings between pairs of entities drawn from each of the two ontologies. A correspondence (a.k.a. a mapping) is defined as a 4-tuple $(e_{1},e_{2},r,c)$, where $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. $r$ is the relationship between two matched entities $e_{1}$ and $e_{2}$, and $c\in[0,1]$ is the confidence of the correspondence (or may be absent). The relationship $r$ in OM tasks can be equivalence ($\equiv$), subsumption ($\subseteq$), or other more complex relationships [2]. An alignment (A) is a set of candidate correspondences, while a reference (R) is a set of gold-standard correspondences verified by domain experts (i.e. the ground truth alignment).
A classical OM system usually contains syntactic matching, lexical matching, and semantic matching. This multilayer architecture has been implemented in several successful OM systems, including LogMap [8, 9], AgreementMakerLight (AML) [10, 11], FCA-Map [12, 13], and Agent-OM [14]. Syntactic matching is often the first step of classical OM. Text preprocessing is commonly used for syntactic matching. It has proven to correctly match many ontology entities because entities with similar string combinations are likely to be correct, and it is computationally efficient.
III Related Work
III-A Text preprocessing in OM
The use of text preprocessing in OM can be traced back to the early stages of OM systems, which were initially developed and widely used as a basic component to generate linguistic-based mappings. AgreementMaker [15] has a normaliser to unify the textual information of the entities. SAMBO [16] uses the Porter Stemmer for each word to improve the similarity measure for terms with different prefixes and suffixes. In RiMOM [17], the context information of each entity is viewed as a document. The text in each document is preprocessed with tokenisation, stop words removal, and stemming. Recently, machine learning (ML) models have emerged for modern OM systems. While text preprocessing remains useful, its role is more focused on normalising the text that becomes the input to the model. For example, BERTMap [18] uses BERT’s inherent WordPiece tokeniser to build the subword of each entity. DeepOnto [19] extends normalisation to axioms using EL embedding models [20]. The ML extension of LogMap [21] reuses the seed mappings of traditional systems, where each entity is split into its component words and the mapping is based on pairwise component similarities.
To the best of our knowledge, most of the literature implements a preprocessing method without explaining why a specific method is chosen, and no studies have evaluated the effect of the entire text preprocessing pipeline.
III-B LLMs in OM
Advances in LLM research raises the question of whether text preprocessing is obsolete for OM. Early explorations in [22, 23] show the potential of using LLMs for OM, but matching by LLM prompting alone commonly results in poor performance. OLaLa [24] proposes a prototype and states that the LLMs for OM include a number of decisions on model and parameter combinations, and there is no one-model-fits-all solution. Works on both Agent-OM [14] and LLM4OM [25] point out that LLMs have limited context length and inherited hallucinations. While LLM4OM highlights the importance of integrating the Retrieval-Augmented Generation (RAG) [26], Agent-OM introduces an agent-based framework and enhances the compatibility of LLMs via function calling [27] for OM to improve matching performance. The authors in [28] state that the use of LLMs for OM has a high computational overhead and apply a prioritised depth-first search (PDFS) strategy on the retrieve-then-prompt pipeline to optimise candidate selection. Recent works in [29, 30, 31] also utilise LLMs and LLM embeddings for complex OM.
However, these works focus on building a sophisticated framework for using LLMs for OM, and less attention has been paid to the subtask of using LLMs for text preprocessing as used in OM. The challenges are twofold. On one hand, LLM hallucinations may produce incorrect “seed mappings”. Such mappings are toxic for the subsequent lexical and semantic matching. On the other hand, it is difficult to select a generic LLM prompting strategy from a number of LLM prompting templates and strategies. With strong background knowledge, LLMs can be helpful in detecting false mappings after text preprocessing. We discuss LLM-based repair in Section V-D.
IV Text Preprocessing Pipeline Analysis
IV-A Experiment Setup
@@
Fig. 2 shows the experiment setup to analyse the text preprocessing pipeline. Alignment (A) is generated via the text preprocessing pipeline. For both $O_{s}$ and $O_{t}$, we retrieve the entities from classes (owl:Class) and properties (both owl:ObjectProperty and owl:DatatypeProperty). For those ontologies where the names of the concepts are not textual (e.g. a code or numerical identifier), we retrieve the meaningful text from entity labels (rdfs:label or other annotation properties) instead. Then, we apply the text preprocessing pipeline method $f(\cdot)$ on each entity $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$. If $f(e_{1})=f(e_{2})$, we store the correspondence in the corresponding alignment file. While OM systems often follow text preprocessing with a more sophisticated strategy to determine the equivalence of entities, we define equivalence to mean identical text generated by preprocessing alone. This enables us to directly evaluate the preprocessing performance as a standalone contribution to OM. We compare the generated alignment (A) with the reference (R) to evaluate the performance of the text preprocessing pipeline on OM.
@@
Figure 2: Experiment setup to analyse the effect of text preprocessing pipeline in OM: Tokenisation (T), Normalisation (N), Stop Words Removal (R), and Stemming/Lemmatisation (S/L).
@@
We select 8 OAEI tracks stored in the Ontology Matching Evaluation Toolkit (MELT) [32] public repository (retrieved January 1, 2025). Table I shows the details of the selected tracks. Each track may contain more than one alignment corresponding to different pairs of ontologies. For example, the Largebio Track has multiple reference files, pairing FMA [33] and NCI [34], FMA and SNOMED [35], and SNOMED and NCI, respectively. The number of alignments for each track is given in the table. There are 49 distinct alignments evaluated in this study. Fig. 3 shows the number of entities in each alignment across different tracks.
TABLE I: Selected OAEI tracks and the number of alignments.
@@
Name Domain # Alignments Anatomy Human and Mouse Anatomy 1 Biodiv Biodiversity and Ecology 9 CommonKG Common Knowledge Graphs 3 Conference Conference 24 Food Food Nutritional Composition 1 Largebio Biomedical 6 MSE Materials Science & Engineering 3 Phenotype Disease and Phenotype 2
@@
Figure 3: Number of entities in each alignment across different tracks.
@@
Compound words are frequently used in ontology naming conventions. For example, compound words such as “art gallery” can be formatted as the Pascal case “ArtGallery” used in YAGO [36] or the Snake case “art_gallery” used in Wikidata [37]. Fig. 4 shows the frequency distribution of compound words in the selected OAEI tracks. For comparisons between entities with compound words, one approach is to use their tokens, where “XY” is equivalent to “YX” because they share the same concepts X and Y. However, our comparison considers both concepts and their order, so that “XY” is not equivalent to “YX”. Although concurrent “XY” and “YX” are rare in the matching process, we assume that it could happen and so we take account of the order of concepts in compound word names to ensure a fair comparison in our experiments.
(a) Classes
(b) Properties
@@
Figure 4: Frequency distribution of compound words. We exclude entities with more than 15 compound words because their proportion is less than 1%.
@@
There are a variety of methods that can be employed in a general text preprocessing pipeline, but not all of them are applicable to syntactic OM. We select the following methods in each text preprocessing step: (a) Tokenisation (T): We use the conventional full word tokenisation method to separate the compound words in an entity name. Experiments on several alternative subword tokenisation methods (e.g. Byte Pair Encoding [38, 39], WordPiece [40, 41], and Unigram [42, 43]) show that results after concatenating the processed name parts back together are generally the same as for full word tokenisation. We do not use sentence tokenisation (a.k.a. segmentation) because text retrieved from the entity’s name or label is commonly short. (b) Normalisation (N): includes lowercasing, HTML tag removal, separator formatting, and punctuation removal. Other methods that may potentially change the word semantics are excluded (e.g. removal of special characters and numbers). (c) Stop Words Removal (R): includes the most common English stop words defined in the Natural Language Toolkit (NLTK) [44]. (d) Stemming/Lemmatisation (S/L): Stemming methods include Porter, Snowball, and Lancaster. Lemmatisation utilises the NLTK Lemmatiser based on WordNet [45], and word categorisation employs POS Tagging.
@@
There are four primitive measures in information retrieval: true positive (TP), false positive (FP), false negative (FN), and true negative (TN). In the context of OM, the evaluation compares an alignment (A) returned by the OM system with a gold standard reference (R) validated by domain experts. Fig. 5 illustrates that the four primitive measures in OM can be interpreted as $TP=A\cap R$, FP=ARFP=A-R, FN=RAFN=R-A, and $TN=(C\times C^{\prime})-(A\cup R)$, where $C\times C^{\prime}$ refers to all possible correspondences. Accuracy (Acc), Specificity (Spec), Precision (Prec), Recall (Rec), and $F_{\beta}$ Score ($F_{\beta}$) are the most common evaluation measures based on TP, FP, FN, and TN. In the context of OM, since $C\times C^{\prime}$ is extremely large (the Cartesian product of $e_{1}\in O_{s}$ and $e_{2}\in O_{t}$), TN is often much larger than TP, FP, and FN. This means that Accuracy (Acc) and Specificity (Spec) are close to 1, and they have no statistically significant difference across different alignments. We note that Precision (Prec) and Recall (Rec) contribute equally to $F_{\beta}$. Therefore, we choose Precision (Prec), Recall (Rec), and F1 Score ($\beta=1$) in this study. They are defined as follows.
Figure 5: OM evaluation [46].